solve by elimination
3k + 7m = 16
4k - 3m = 9

Answer by mananth(16949)

3k+7m=16 .............1
4k-3m=9 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 7
9k+21m= 48
28k-21m =63
Add the two equations
37k= 111
/37
k=3
plug value of k in (1)
3k+7m=16
9+7m=16
7m=16-9
7m=7
m=1